Two cyclists start biking from a trail's start 3 hours apart. The second cyclist travels at 10 miles per hour and starts 3 hours after the first cyclist who is traveling at 6 miles per hour. How much time will pass before the second cyclist catches up with the first from the time the second cyclist started biking?
Answer by dhartman0220(8)

Distance is equal to velocity multiplied by time D=v*t
The distance traveled by the second biker will equal the distance traveled by the first at some point in time.
hence D1=first cyclist
v=6(m/h)
t=(3+x)h
D2=second cyclist
v=10(m/h)
t=xh
we know D1 is 3 hours ahead and D2 has to travel (x) amount of time to catch up.
so
D1=D2
6(3+x)=10x
18+6x=10x
18=4x
x=4.5 hours